OpenStudy (anonymous):

PLEASE HELP! The claim is that the proportion of peas with yellow pods is equal to 0.25 (or 25%). The sample statistics from one experiment include 570 peas with 156 of them of them having yellow pods. Find the value of the test statistic.

OpenStudy (anonymous):

That test statistic for a proportion is \[Z=\frac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}\] where \(\hat{p}=\dfrac{156}{570}\approx0.274\) is the sample proportion, \(p_0=0.25\) is the proportion assumed under the null hypothesis, and \(n=570\) is the sample size.

OpenStudy (anonymous):

So how would I obtain the value size?

OpenStudy (anonymous):

All you have to do is plug in the right numbers and compute. \[\frac{0.274-0.25}{\sqrt{\dfrac{0.25(1-0.25)}{570}}}\] http://www.wolframalpha.com/input/?i=%280.274-0.25%29%2FSqrt%5B0.25%281-0.25%29%2F570%5D